# A Cramér–Wold theorem for mixtures

### v1

@@ 5. Concluding remarks
Gaussian-mixture models have been shown to be very effective in modeling different real data, see for instance Titterington et al. [19] for a deep study of their properties. They provide flexible and general models, and relevant applications can be found in the literature in different fields like density estimation, machine learning and clustering, among others. The estimation of these models is quite involved, in particular for high-dimensional data, typically using Markov-chain Monte-Carlo methods in a Bayesian framework. We believe that the Cramér–Wold device may well have a role to play in this circle of ideas, in particular through Theorems 3.1 and 3.2. We intend to explore this in a future paper.

### v2

@@ 5. Concluding remarks
Gaussian-mixture models have been shown to be very effective in modeling different real data, see for instance Titterington et al. [19] for a deep study of their properties. They provide flexible and general models, and relevant applications can be found in the literature in different fields like density estimation, machine learning and clustering, among others. The estimation of these models is quite involved, in particular for high-dimensional data, typically using Markov-chain Monte-Carlo methods in a Bayesian framework. The Cramér–Wold device has a role to play in this circle of ideas, in particular through Theorems 4.1 and 4.2. This is explored in detail in the companion paper [8].